[figure: 9.7 cm]

DESTRUCTION OF ANTIFERROMAGNETIC ORDER BY LOCALIZED POLARONS

Z. Wilamowski and E. Hankiewicz Institute of Physics, Polish Academy of Sciences, Warsaw, Poland

The formation of Mott polarons is connected with the breaking of antiferromagnetic order and the induction of magnetic moments by strong coupling. Here we show that also weaker p-d coupling suitable for HT_{C}-superconductors may lead to localization of carriers and reduction of the AF correlation range. We consider a 2D Heisenberg antiferromagnet composed from localized d-spins which interacts with p hole by exchange coupling and solve the problem exactly. We find that p-d interaction lowers the energy inversely proportional to the polaron localization range σ. Since the kinetic energy increases as a square of 1/σ the total energy has a minimum. The corresponding localization radius σ is of the order of 10 Å for CuO_{2} planes. We show that in our approach localization leads to a reduction of AF correlation range by compensation and adjustment of AF spins to a hole perturbation. Magnetic properties of the localized polarons, and the resulting destruction of the long range antiferromagnetic order correspond well to experimental results.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Zbysław Wilamowski

Address for correspondence : Instytut Fizyki PAN al. Lotników 32/46 02–668 Warszawa

Email address : wilamz@ifpan.edu.pl